(x^2-(9x+5)/(x^2+1x+6) + (3x^2-6-6x)/(-6-1x-x^2)=A/B
A=
B=

Answer by Eezzdboy(5)

A=3/x-1 and B=-1+or -sqrt2/x^2 -1